What is the tenth term of the sequence that has a common ratio of 1/3 and 36 as it's fifth term?
Answer by fcabanski(1391)

a(n)=a(x)r^(x-5) where x is the number of the known term, in this case 5th (5), and n is the number of the term to be found, in this case 10th (10), and r is the common ratio.


a(10)=36*(1/3)^(10-5) = 36*(1/3)^5 = .148148...